# R1-VL: Learning to Reason with Multimodal Large Language Models via Step-wise Group Relative Policy Optimization

### v1

@@ Abstract
Recent studies generally enhance MLLMs’ reasoning capabilities via supervised fine-tuning on high-quality chain-of-thought reasoning data, which often leads models to merely imitate successful reasoning paths without understanding what the wrong reasoning paths are. In this work, we aim to enhance the MLLMs’ reasoning ability beyond passively imitating positive reasoning paths. To this end, we design Step-wise Group Relative Policy Optimization (StepGRPO), a new online reinforcement learning framework that enables MLLMs to self-improve reasoning ability via simple, effective and dense step-wise rewarding. Specifically, StepGRPO introduces two novel rule-based reasoning rewards: Step-wise Reasoning Accuracy Reward (StepRAR) and Step-wise Reasoning Validity Reward (StepRVR). StepRAR rewards the reasoning paths that contain necessary intermediate reasoning steps via a soft key-step matching technique, while StepRAR rewards reasoning paths that follow a well-structured and logically consistent reasoning process through a reasoning completeness and logic evaluation strategy. With the proposed StepGRPO, we introduce R1-VL, a series of MLLMs with outstanding capabilities in step-by-step reasoning. Extensive experiments over 8 benchmarks demonstrate the superiority of our methods. Code will be available at link.
@@ 2.1 Multimodal Large Language Model
Multimodal Large Language Models (MLLMs) [13, 1, 34, 46, 16, 7, 40, 19, 47] have shown remarkable advancements across a wide range of vision-language understanding tasks, demonstrating their capabilities in comprehending and analyzing visual contents across various application domains. Early research on MLLMs primally focuses on text generation based on text prompts and input multiple modalities such as images [19, 18], videos [31, 8]. Recent advancements further enhance the capabilities of MLLMs from various aspects. For example, recent models [23, 39] incorporate multimodal inputs and outputs such as video, audio, and point cloud inputs beyond text and images. In addition, some efforts attempt to adapt MLLMs for domain-specific tasks, such as medical image understanding [50, 17, 15] and document analysis [44, 20]. In this work, we focus on enhancing the reasoning ability of MLLMs in tackling complex reasoning tasks and introduce R1-VL, a series of MLLMs that have superior reasoning capability.
@@ 2.2 MLLM Reasoning
Inspired by the advances in NLP that show great potential in learning to reason and tackling complex language tasks [26], recent studies attempt to enhance the reasoning capability of MLLM. Generally, current MLLM reasoning methods improve the reasoning capability of MLLM by generating high-quality chain-of-thoughts (CoT) data using powerful model (e.g., GPT-4) and performing supervised fine-tuning with the collected data [41, 49, 33, 42, 9]. For example, LLaVA-COT [41] prompts GPT-4 to generate CoT data with structured reasoning templates, i.e., summary, caption, reasoning and conclusion, and then performs SFT on the generated data. Mulberry [42] introduces Collective Monte Carlo Tree Search (MCTS) into MLLM and proposes CoMCTS which leverages complementary knowledge from multiple models to collaboratively search and identify effective reasoning paths. Different from these works, we aim to enhance the MLLMs’ reasoning ability via self-exploration through online reinforcement learning and design StepGRPO that enables MLLM to self-improve the reasoning ability with step-wise reward signals.
@@ 2.3 Reinforcement Learning
Reinforcement Learning (RL) [14] is a fundamental approach in machine learning, where an agent learns to interact with an environment by taking actions, receiving rewards, and updating its policy to maximize the long-term return. Typical RL methods, such as Q-learning [38], have been widely applied in robotics, game playing (e.g., AlphaGo), and autonomous control. With the rise of large language models (LLMs) [27, 3, 25], Reinforcement Learning with Human Feedback (RLHF) [2] has emerged as a key technique for fine-tuning models using human preference data. RLHF leverages algorithms like Proximal Policy Optimization (PPO) [29] and Direct Preference Optimization (DPO) [28] to guide model behavior for improving the alignment, coherence and helpfulness in response generation.
@@ 3.2.2 Step-wise Online Policy Optimization
Step-wise reasoning accuracy reward (StepRAR) reduces the effect of learning from sparse reward by additionally rewarding reasoning paths that contain correct intermediate reasoning steps contributing to the final solution. Specifically, for each question $Q$, we pre-extract a set of key reasoning steps $\mathbf{v}=\{v_{1},v_{2},...\}$ from the corresponding reasoning path $\tau$ in dataset $D_{s}$. We define key steps as the essential variables and equations that directly contribute to the final solution, and prompt GPT-4 to extract several key steps from the reasoning path for each question.
@@ 4 Experiment
Method MathVista MMStar Math-V ChartQA DynaMath HallBench MathVerse MMEsum AVG
@@
Cambrain-1-8B [34] 49.0 - - 73.3 - - - - -
MM-1.5-7B [46] 47.6 - - 78.6 - - - 1861 -
Idefics3-LLaMA3-8B [16] 58.4 55.9 - 74.8 - - - 1937 -
InternVL2-8B [7] 58.3 61.5 - 83.3 39.7 - - 2210 -
MiniCPM-V-2.6-8B [43] 60.6 57.5 - - - 48.1 - 2348 -
DeepSeek-VL2-MOE-4.5B [40] 62.8 61.3 - 86.0 - - - 2253 -
@@
LLaVA-CoT-11B [41] 54.8 57.6 - - - 47.8 - - -
LLaVA-Reasoner-8B [49] 50.6 54.0 - 83.0 - - - - -
Insight-V-8B [9] 49.8 57.4 - 77.4 - - - 2069 -
Mulberry-7B [42] 63.1 61.3 - 83.9 45.1 54.1 - 2396 -
@@
Qwen2-VL-2B [37] 43.0 48.0 12.4 73.5 24.9 41.7 19.7 1872 41.2
Qwen2-VL-2B-GRPO [30] 41.4 46.2 16.0 72.5 24.2 42.2 19.9 1930 41.4
R1-VL-2B (Ours) 52.1 49.8 17.1 75.2 29.4 44.0 26.2 2048 45.8
Qwen2-VL-7B [37] 58.2 60.7 16.3 83.0 42.1 50.6 32.5 2327 53.3
Qwen2-VL-7B-GRPO [30] 55.1 59.8 19.1 81.3 33.9 48.5 30.9 2335 51.4
R1-VL-7B (Ours) 63.5 60.0 24.7 83.9 45.2 54.7 40.0 2376 57.1
Table 1: Main experimental results. To comprehensively examine the proposed StepGRPO, we conduct extensive experiments with two baseline models on eight benchmarks, and compare StepGRPO with various state-of-the-art MLLMs.
@@ 4.1 Datasets
For policy warm-up, we adopt Mulberry-260k [42] for supervised fine-tuning. For step-wise online policy optimization, we randomly sample 10K data from Mulberry-260k as our training data. For evaluation, we adopt 8 widely-used multimodal benchmarks for comprehensively evaluating our proposed StepGRPO, including MathVista [21], MMStar [5], Math-Vision [36], ChartQA [24], DynaMath [51], HallusionBench [11], MathVerse [48] and MME [10]. These multimodal benchmarks cover a wide range of tasks from mathematical reasoning, chart understanding, visual hallucination and general visual understanding.
@@ 4.3 Main Experimental Results
We first compare R1-VL with its baseline models, Qwen2-VL-2B and Qwen2-VL-7B. From Table 1, we observe that applying GRPO directly to baseline models often results in performance degradation, primarily due to the sparse reward issue. The baseline models exhibit limited reasoning capability, leading to very few reasoning paths receiving rewards, which negatively impacts the reasoning capability. In contrast, R1-VL with our proposed StepGRPO consistently improves the baseline models by significant margins, achieving 4.6% improvement over Qwen2-VL-2B and 3.8% over Qwen2-VL-7B. This improvement is largely attributed to that StepGRPO introduces step-wise reasoning accuracy and validity rewards, which provide rich and informative supervision at each reasoning step, effectively mitigating the sparse reward issue for MLLMs.

### v2

@@ Abstract
Recent studies generally enhance MLLMs’ reasoning capabilities via supervised fine-tuning on high-quality chain-of-thought reasoning data, which often leads models to merely imitate successful reasoning paths without understanding what the wrong reasoning paths are. In this work, we aim to enhance the MLLMs’ reasoning ability beyond passively imitating positive reasoning paths. To this end, we design Step-wise Group Relative Policy Optimization (StepGRPO), a new online reinforcement learning framework that enables MLLMs to self-improve reasoning ability via simple, effective and dense step-wise rewarding. Specifically, StepGRPO introduces two novel rule-based reasoning rewards: Step-wise Reasoning Accuracy Reward (StepRAR) and Step-wise Reasoning Validity Reward (StepRVR). StepRAR rewards the reasoning paths that contain necessary intermediate reasoning steps via a soft key-step matching technique, while StepRAR rewards reasoning paths that follow a well-structured and logically consistent reasoning process through a reasoning completeness and logic evaluation strategy. With the proposed StepGRPO, we introduce R1-VL, a series of MLLMs with outstanding capabilities in step-by-step reasoning. Extensive experiments over 8 benchmarks demonstrate the superiority of our methods. Code is available at link.
@@ 2.1 Multimodal Large Language Model
Multimodal Large Language Models (MLLMs) [15, 1, 38, 51, 18, 8, 43, 21, 52] have shown remarkable advancements across a wide range of vision-language understanding tasks, demonstrating their capabilities in comprehending and analyzing visual contents across various application domains. Early research on MLLMs primally focuses on text generation based on text prompts and input multiple modalities such as images [21, 20, 53], videos [35, 9]. Recent advancements further enhance the capabilities of MLLMs from various aspects. For example, recent models [25, 42] incorporate multimodal inputs and outputs such as video, audio, and point cloud inputs beyond text and images. In addition, some efforts attempt to adapt MLLMs for domain-specific tasks, such as medical image understanding [56, 19, 17] and document analysis [49, 22]. In this work, we focus on enhancing the reasoning ability of MLLMs in tackling complex reasoning tasks and introduce R1-VL, a series of MLLMs that have superior reasoning capability.
@@ 2.2 MLLM Reasoning
Inspired by the advances in NLP that show great potential in learning to reason and tackling complex language tasks [29], recent studies attempt to enhance the reasoning capability of MLLM. Generally, current MLLM reasoning methods improve the reasoning capability of MLLM by generating high-quality chain-of-thoughts (CoT) data using powerful model (e.g., GPT-4) and performing supervised fine-tuning with the collected data [44, 55, 37, 46, 10]. For example, Mulberry [46] introduces Collective Monte Carlo Tree Search (MCTS) into MLLM and proposes CoMCTS which leverages complementary knowledge from multiple models to collaboratively search and identify effective reasoning paths. In addition, recent works [47, 30, 27, 14] attempt to explore online reinforcement learning to improve the MLLMs’ reasoning ability. Different from these works, we design StepGRPO that enables MLLM to self-improve the reasoning ability with step-wise reward signals.
@@ 2.3 Reinforcement Learning
Reinforcement Learning (RL) [16] is a fundamental approach in machine learning, where an agent learns to interact with an environment by taking actions, receiving rewards, and updating its policy to maximize the long-term return. With the rise of large language models (LLMs) [31, 4, 28], Reinforcement Learning with Human Feedback (RLHF) [3] has emerged as a key technique for fine-tuning models using human preference data. RLHF leverages algorithms like Proximal Policy Optimization (PPO) [33] and Direct Preference Optimization (DPO) [32] to guide model behavior for improving the alignment, coherence and helpfulness in response generation.
@@ 4 Experiment
Method MathVista MMStar Math-V ChartQA DynaMath HallBench MathVerse MMEsum MMReason
AVG
@@
Cambrain-1-8B [38] 49.0 - - 73.3 - - - -
MM-1.5-7B [51] 47.6 - - 78.6 - - - 1861
Idefics3-LLaMA3-8B [18] 58.4 55.9 - 74.8 - - - 1937
InternVL2-8B [8] 58.3 61.5 - 83.3 39.7 - - 2210
MiniCPM-V-2.6-8B [48] 60.6 57.5 - - - 48.1 - 2348
DeepSeek-VL2-MOE-4.5B [43] 62.8 61.3 - 86.0 - - - 2253 11.5
@@
LLaVA-CoT-11B [44] 54.8 57.6 - - - 47.8 - -
LLaVA-Reasoner-8B [55] 50.6 54.0 - 83.0 - - - -
Insight-V-8B [10] 49.8 57.4 - 77.4 - - - 2069
Mulberry-7B [46] 63.1 61.3 - 83.9 45.1 54.1 - 2396 11.8
@@
Vision-R1-7B [14] 73.5 - - - - - 52.4 - -
LMM-R1 [30] 63.2 58.0 26.3 - - - 41.5 - -
R1-ShareVL-7B [47] 75.4 67.0 29.5 - - - 52.8 - -
Qwen2-VL-2B [41] 43.0 48.0 12.4 73.5 24.9 41.7 19.7 1872 7.7 37.5
R1-VL-2B (Ours) 52.1 49.8 17.1 75.2 29.4 44.0 26.2 2048 8.3 41.6
Qwen2-VL-7B [41] 58.2 60.7 16.3 83.0 42.1 50.6 32.5 2327 11.9 48.7
R1-VL-7B (Ours) 63.5 60.0 24.7 83.9 45.2 54.7 40.0 2376 12.5 52.1
Qwen2.5-VL-7B [2] 68.2 63.9 25.1 87.3 53.2 52.1 49.2 2347 17.3 55.5
R1-VL-7B* (Ours) 74.3 66.2 28.2 87.7 56.5 57.2 52.2 2395 17.9 58.4
Table 1: Main experimental results. To comprehensively examine the proposed StepGRPO, we conduct extensive experiments with two baseline models on eight benchmarks, and compare StepGRPO with various state-of-the-art MLLMs.* indicates that the model is trained using Qwen2.5-VL-7B as the base model with the data from [47].
@@ 4.1 Datasets
For policy warm-up, we adopt Mulberry-260k [46] for supervised fine-tuning. For step-wise online policy optimization, we randomly sample 10K data from Mulberry-260k as our training data. For evaluation, we adopt 8 widely-used multimodal benchmarks for comprehensively evaluating our proposed StepGRPO, including MathVista [23], MMStar [6], Math-Vision [40], ChartQA [26], DynaMath [57], HallusionBench [12], MathVerse [54], MME [11] and MMReason [45]. These multimodal benchmarks cover a wide range of tasks from mathematical reasoning, chart understanding, visual hallucination and general visual understanding.
@@ 4.3 Main Experimental Results
We first compare R1-VL with its baseline models, Qwen2-VL-2B and Qwen2-VL-7B. The baseline models exhibit limited reasoning capability, leading to very few reasoning paths receiving rewards, which negatively impacts the reasoning capability. In contrast, R1-VL with our proposed StepGRPO consistently improves the baseline models by significant margins, achieving 4.6% improvement over Qwen2-VL-2B and 3.8% over Qwen2-VL-7B. This improvement is largely attributed to that StepGRPO introduces step-wise reasoning accuracy and validity rewards, which provide rich and informative supervision at each reasoning step, effectively mitigating the sparse reward issue for MLLMs.
@@ 4.5 Discussion
We provide more discussions, experimental results and qualitative analysis in the appendix.
@@ 5 Conclusion
Acknowledgement. This research is supported by the RIE2025 Industry Alignment Fund – Industry Collaboration Projects (IAF-ICP) (Award I2301E0026), administered by A*STAR, as well as supported by Alibaba Group and NTU Singapore through Alibaba-NTU Global e-Sustainability CorpLab (ANGEL).